# AI-Face: A Million-Scale Demographically Annotated AI-Generated Face Dataset and Fairness Benchmark

### v1

@@ 4 Fairness Benchmark Experiments
Figure 5: Ratios of FPR for each subgroup to reference subgroup Male-White (M-W). The blue line indicates the 50% margin above. Red and green bars indicate the above and below of the margin, respectively.
Fairness Robustness Evaluation. Images spread on public platforms usually undergo post-processing. Therefore, it is important to estimate the capability of detectors to preserve fairness robustness while handling distorted images. We apply 6 post-processing methods: Random Crop (RC) [76], Rotation (RT) [25], Brightness Contrast (BC) [25], Hue Saturation Value (HSV) [25], Gaussian Blur (GB) [25], and JEPG Compression (JC) [77] to the test images (see Appendix B.5 for more details). Fig. 6 shows each detector’s intersectional $F_{EO}$ and AUC performance changes after using post-processing. Our observations are: 1) These impairments tend to wash out forensic traces, to the point that detectors have significant performance degradation. 2) Recent Fairness-enhanced detectors struggle to maintain fairness when images undergo post-processing. 3) Transform-based models (i.e., ViT-B/16 [63] and UnivFD [67]) demonstrate stronger robustness compared with CNN-based models. 4) JEPG Compression and Gaussian Blur cause notably greater performance degradation compared to others. See Appendix B.6 for more robustness analysis with respect to different degrees of post-processing.
Figure 6: Visualization of performance changes after post-processing. Shorter bar represents better robustness.
@@
Model Type Detector $F_{OAE}$ $F_{EO}$ AUC $F_{OAE}$ $F_{EO}$ AUC $F_{OAE}$ $F_{EO}$ AUC Avg-$F_{R}$
Xception [61] 4.227(+3.956) 9.198(+8.759) 82.479(-17.289) 2.308(+2.037) 8.691(+8.252) 75.933(-23.835) 0.438(+0.167) 1.724(+1.285) 94.315(-5.453) 5.167
EfficientB4 [62] 3.689(+3.094) 17.017(+15.788) 61.436(-37.681) 4.459(+3.864) 10.191(+8.962) 63.871(-35.246) 0.001(-0.594) 3.621(+2.392) 87.522(-11.595) 8.000
Naive ViT-B/16 [63] 4.45(+4.028) 9.154(+8.919) 70.896(-28.018) 2.531(+2.109) 5.557(+5.322) 68.935(-29.979) 1.249(+0.827) 2.874(+2.639) 89.109(-9.805) 6.667
F3Net [64] 1.749(+1.663) 19.484(+18.932) 86.265(-13.561) 2.995(+2.909) 5.445(+4.893) 82.421(-17.405) 0.155(+0.069) 2.927(+2.375) 93.882(-5.944) 6.000
SPSL [65] 8.497(+8.309) 2.430(+1.853) 75.177(-24.609) 3.323(+3.135) 8.966(+8.389) 82.024(-17.762) 0.138(-0.050) 2.321(+1.744) 94.320(-5.466) 6.167
Frequency SRM [66] 3.708(+3.440) 1.169(+0.633) 65.779(-33.988) 4.976(+4.708) 33.702(+33.166) 72.777(-26.990) 1.545(+1.277) 2.378(+1.842) 94.130(-5.637) 8.000
UCF [16] 2.930(+2.761) 9.924(+9.578) 83.260(-16.625) 3.536(+3.367) 9.395(+9.049) 83.92(-15.965) 1.346(+1.177) 1.377(+1.031) 94.948(-4.937) 6.500
UnivFD [67] 14.149(+13.592) 1.833(+1.343) 65.810(-30.220) 7.686(+7.129) 11.701(+11.211) 69.483(-26.547) 0.903(+0.346) 2.227(+1.737) 85.965(-10.065) 8.167
Spatial CORE [68] 0.308(-0.669) 11.854(+10.008) 79.222(-19.624) 3.966(+2.989) 5.267(+3.421) 81.264(-17.582) 0.005(-0.972) 2.943(+1.097) 94.329(-4.517) 5.667
DAW-FDD [20] 5.040(+4.917) 4.993(+4.294) 80.308(-19.395) 2.577(+2.454) 7.253(+6.554) 78.562(-21.141) 0.205(+0.082) 2.708(+2.009) 93.876(-5.827) 6.000
DAG-FDD [20] 4.279(+4.087) 13.565(+13.158) 85.859(-14.012) 3.885(+3.693) 7.350(+6.943) 83.153(-16.718) 1.062(+0.870) 1.688(+1.281) 94.326(-5.545) 7.167
Fairness- enhanced PG-FDD [21] 4.263(+4.129) 11.077(+10.840) 81.174(-18.763) 1.984(+1.850) 4.715(+4.478) 84.572(-15.365) 1.205(+1.071) 1.159(+0.922) 94.962(-4.975) 4.500
@@
Table 6: Fairness generalization results based on the gender attribute. The smallest performance changes (in parentheses) and the best performance are in bold and in red, respectively.
Fairness Generalization Evaluation. To evaluate detectors’ fairness generalization capability, we train them on AI-Face and test them on A-DF-1.0, DF-Platter, and GenData, none of which are part of AI-Face. Results on gender attribute in Table 6 show that: 1) According to Avg-$F_{R}$, the top three methods excelling in fairness preservation are PG-FDD, Xception, and CORE. PG-FDD, specifically designed for fairness generalization, leads to overall performance. However, it does not excel in terms of performance changes compared with intra-domain test results from Table 5, indicating room for improvement in its generalization capabilities. 2) CORE is notable for demonstrating negative fairness performance changes on A-DF-1.0 and GenData, suggesting techniques within CORE that could be potentially explored to enhance fairness generalization. More results are in Appendix B.7.
@@
Figure 7: Impact of training set size on detectors’ intersectional $F_{EO}$(%) and AUC (%).
Effect of Increasing Training Set Size. We randomly sample 20%, 40%, 60%, and 80% of each training subset from AI-Face to assess the impact of training size on performance. Key observations from Fig. 7: 1) The performance of UnivFD changes slightest and cannot be improved with the increasing of data size. 2) Overall, detectors’ performance improves with larger training size, though few show fluctuations (e.g., ViT-B/16 and CORE). 3) A larger training set may improve utility but not always fairness. For example, Xception and SRM show increased utility when training size grows from 60% to 80%, but fairness worsens. Similar trends are observed in DAG-FDD and SPSL when the training set size increases from 40% to 60%. See Appendix B.8 for full results.
@@
Discussion. According to the above experiments, we summarize the unsolved fairness problems in recent detectors: 1) Detectors’ fairness is unstable when detecting face images generated by different generative methods, indicating a future direction for enhancing fairness stability since new generative models continue to emerge. 2) Even though fairness-enhanced detectors exhibit small overall fairness metrics, they still show biased detection towards minority groups. Future studies should be more cautious when designing fair detectors to ensure balanced performance across all demographic groups. 3) There is currently no reliable detector, as all detectors experience severe large performance degradation under image post-processing and cross-domain evaluation. Future studies should aim to develop a unified framework that ensures fairness, robustness, and generalization, as these three characteristics are essential for creating a reliable detector.
@@ 5 Conclusion
This work presents the first demographically annotated million-scale AI-Face dataset, serving as a pivotal foundation for addressing the urgent need for developing fair AI face detectors. Based on our AI-Face dataset, we conduct the first comprehensive fairness benchmark, shedding light on the fairness performance and challenges of current representative AI face detectors. Our findings can inspire and guide researchers in refining current models and exploring new methods to mitigate bias. Limitation and Future Work: One limitation is that age annotations in our AI-Face dataset have relatively lower accuracy as the age attribute is often too ambiguous to predict. We will improve our annotator’s accuracy in predicting age attributes in the future. Additionally, we plan to extend our fairness benchmark to evaluate large language models like LLaMA2 [78] and GPT4 [79] for detecting AI faces. Social Impact: Malicious users could misuse AI-generated face images from our dataset to create fake social media profiles and spread misinformation. To mitigate this risk, only users who submit a signed end-user license agreement (EULA) will be granted access to our dataset.
Acknowledgment
This work is supported by the U.S. National Science Foundation (NSF) under grant IIS-2348419 and the National Artificial Intelligence Research Resource (NAIRR) Pilot and TACC Lonestar6. The views, opinions and/or findings expressed are those of the author and should not be interpreted as representing the official views or policies of NSF and NAIRR Pilot.
@@ Acknowledgment
Appendix A The Details of Demographically Annotated AI-Face Dataset
A.1 Phase1: Annotator Development
A.1.1 Annotator Implementation Details
For developing the annotator, all experiments are based on the PyTorch with a single NVIDIA RTX A6000 GPU. For training, we fix the batch size 64, epochs 32, and use Adam optimizer with an initial learning rate $\beta=1e-3$. Additionally, we employ a Cosine Annealing Learning Rate Scheduler to modulate the learning rate adaptively across the training duration. The hyperparameter $\gamma$ in SAM optimization is set as 0.05. For uncertainty estimation, $k$ and $\rho$ in uncertainty score $V(X^{(a)})$ are set as 100 and 0.2, respectively.
A.1.2 Details of Threshold Settings for Sample Difficulty Level
For Q2, Setting: According to the distribution as shown in Appendix A.2.2, for VGGFace2 [47], A-DFDC [80], and A-DFD [80] test set, the threshold $t_{1}^{Gender}$ and $t_{2}^{Gender}$ are set as 0.25 and 0.4, respectively. And $t_{1}^{Age}$ and $t_{2}^{Age}$ are set as 0.3 and 0.5, respectively. The threshold for gender attribute is more strict than age because gender attribute prediction is a relatively easier task than age, as well as reflecting from the distribution. For A-FF++ [80] and A-Celeb-DF-v2 [80], we adjust the threshold $t_{1}^{Gender}$ to 0.21 and $t_{2}^{Gender}$ to 0.25 in order to get sufficient 1,500 images in each sample difficulty level subset, especially for ‘Hard’ level.
A.1.3 Additional Annotator Evaluation Results
From Table 7 to Table 11 are comparison results of our annotator against baselines InsightFace [43] and Face++[42] on detailed attributes. The findings and results align with the results in Table 3 of the submitted manuscript. For cross-domain evaluation, we additionally choose Adience [81] dataset, where images are manually annotated, consisting of over 26.5k real images of over 2.2k different individuals in unconstrained environments, to further validate the effectiveness and good generalization capability of our annotator. Results in table 12 demonstrate our annotator outperforms InsightFace [43] and Face++[42] again. Overall, one intra-domain dataset (VGGFace2) and five cross-domain datasets (A-FF++, A-DFDC, A-DFD, A-Celeb-DF-v2, and Adience) all validate that our annotator’s superior performance against current state-of-the-art face attribute prediction tools Face++ [42] and InsightFace [43].
Level Method VGGFace2 [47]
Female Male Young Middle_Aged Senior
precision recall F1 precision recall F1 precision recall F1 precision recall F1 precision recall F1
77.042 83.597 80.060 79.571 72.314 75.584 80.201 34.649 46.991 46.049 63.692 53.327 66.266 77.261 71.320
Face++ [42] (0.363) (0.789) (0.444) (0.780) (0.516) (0.453) (2.448) (1.182) (1.278) (0.832) (1.588) (0.976) (1.030) (1.188) (0.955)
76.560 78.062 77.281 76.946 75.395 76.139 78.730 30.533 42.907 41.660 56.653 47.936 61.426 76.107 67.966
InsightFace [43] (0.533) (0.737) (0.534) (0.629) (0.555) (0.474) (2.135) (0.961) (1.052) (1.005) (1.429) (1.078) (0.649) (1.212) (0.724)
81.452 89.467 85.158 87.267 78.329 82.401 95.619 80.467 86.659 91.189 75.027 81.665 88.043 76.720 81.747
All Ours (0.413) (0.331) (0.323) (0.393) (0.572) (0.438) (0.565) (1.343) (0.629) (0.812) (0.880) (0.585) (0.982) (1.388) (0.837)
97.482 96.742 97.108 96.697 97.439 97.064 89.360 57.507 69.964 58.673 72.205 64.729 79.536 89.546 84.240
Face++ [42] (0.604) (0.549) (0.329) (0.545) (0.642) (0.355) (0.670) (1.825) (1.403) (0.837) (1.397) (0.699) (1.119) (0.600) (0.687)
97.625 96.373 96.994 96.420 97.653 97.032 88.544 50.080 63.957 53.146 65.720 58.762 73.657 88.200 80.271
InsightFace [43] (0.403) (0.229) (0.124) (0.206) (0.410) (0.135) (0.917) (1.831) (1.585) (0.621) (1.001) (0.497) (0.808) (0.748) (0.555)
99.575 99.893 99.734 99.893 99.573 99.733 99.720 99.760 99.740 99.279 99.000 99.139 99.551 96.480 97.988
Easy Ours (0.176) (0.100) (0.126) (0.100) (0.177) (0.127) (0.098) (0.080) (0.049) (0.267) (0.400) (0.164) (0.431) (0.688) (0.204)
72.977 81.336 76.927 78.435 69.245 73.549 76.257 24.427 36.996 41.197 61.838 49.446 62.616 73.719 67.710
Face++ [42] (0.292) (1.246) (0.679) (1.101) (0.681) (0.618) (2.732) (0.709) (1.057) (0.961) (2.458) (1.472) (1.171) (0.503) (0.759)
73.710 75.360 74.521 74.807 73.120 73.950 74.222 22.080 34.033 37.372 53.480 43.995 58.070 73.840 65.008
InsightFace [43] (0.699) (1.316) (0.907) (1.070) (0.766) (0.753) (2.488) (0.588) (0.928) (1.208) (2.160) (1.535) (0.282) (1.216) (0.497)
82.518 95.253 88.428 94.389 79.813 86.489 96.648 84.200 89.987 95.382 75.920 84.540 87.960 76.800 81.993
Medium Ours (0.621) (0.482) (0.439) (0.544) (0.858) (0.583) (0.627) (1.730) (0.111) (0.604) (1.017) (0.650) (0.915) (1.544) (1.018)
60.667 72.714 66.146 63.582 50.259 56.140 74.987 22.012 34.013 38.276 57.032 45.807 56.647 68.517 62.009
Face++ [42] (0.194) (0.571) (0.323) (0.694) (0.226) (0.385) (3.942) (1.012) (1.374) (0.697) (0.910) (0.758) (0.799) (2.462) (1.420)
58.345 62.453 60.329 59.611 55.413 57.435 73.425 19.440 30.730 34.462 50.760 41.050 52.550 66.280 58.618
InsightFace [43] (0.498) (0.667) (0.570) (0.610) (0.489) (0.533) (2.999) (0.463) (0.642) (1.187) (1.127) (1.202) (0.858) (1.671) (1.121)
62.263 73.255 67.312 67.518 55.600 60.982 90.490 57.440 70.249 78.905 50.160 61.315 76.617 56.880 65.260
Hard Ours (0.443) (0.410) (0.403) (0.534) (0.680) (0.604) (0.971) (2.218) (1.726) (1.564) (1.222) (0.941) (1.600) (1.933) (1.288)

### v3

@@ 5.1 General Fairness Comparison
Figure 5: FPR(%) of each intersectional subgroup The dashline represents the lowest FPR on Female-Light (F-L) subgroup.
Performance on Different Subsets. 1) Fig. 4 demonstrates the intersectional $F_{EO}$ and AUC performance of detectors on each test subset. We observe that the fairness performance varies a lot among different generative methods for every detector. The largest bias on most detectors comes from detecting face images generated by diffusion models. 2) DAG-FDD [29] and SRM [87] demonstrate the most consistent fairness across subsets, indicating a robust handling of bias introduced by different generative methods. 3) Moreover, the stable utility demonstrates our dataset’s expansiveness and diversity, enabling effective training to detect AI-generated faces from various generative techniques.
Performance on Different Subgroups. We conduct an analysis of all detectors on intersectional subgroups. 1) As shown in Fig. 5, facial images with lighter skin tone are more often misclassified as fake, likely due to the underrepresentation of lighter tones (Tone 1-3) in our dataset (see Fig. 3 (b)). This suggests detectors tend to show higher error rates for minority groups. 2) Although gender representation is relatively balanced (see Fig. 3 (c)) in our dataset, the detectors consistently exhibit higher false positive rates for female subgroups, indicating a persistent gender-based bias.
5.2 Fairness Reliability Assessment
Fairness Robustness Evaluation. We apply 6 post-processing methods: Random Crop (RC) [97], Rotation (RT) [34], Brightness Contrast (BC) [34], Hue Saturation Value (HSV) [34], Gaussian Blur (GB) [34], and JEPG Compression (JC) [98] to the test images. Fig. 6 shows each detector’s intersectional $F_{EO}$ and AUC performance changes after using post-processing. Our observations are: 1) These impairments tend to wash out forensic traces, so that detectors have evident performance degradation. 2) Post-processing does not always cause detectors more bias (e.g., UCF, UnivFD, CORE, DAW-FDD have better fairness after rotation), though they hurt the utility. 3) Fairness-enhanced detectors struggle to maintain fairness when images undergo post-processing. 4) Spatial detectors have better fairness robustness compared with other model types.
Figure 6: Performance ratio after vs. before post-processing. Points closer to 1.0 (i.e., no post-processing) indicate better robustness.
@@ 5.2 Fairness Reliability Assessment
Model Type Detector $F_{OAE}$ ACC $F_{OAE}$ $F_{EO}$ AUC $F_{OAE}$ $F_{EO}$ AUC
Xception 1.006(+0.031) 86.465(-9.843) 6.836(+5.861) 9.789(+5.646) 81.273(-17.310) 2.539(+1.564) 13.487(+9.344) 96.971(-1.612)
EfficientB4 4.077(+0.259) 82.980(-11.223) 8.786(+7.299) 12.370(+6.507) 67.694(-30.917) 3.304(+1.817) 1.995(-3.686) 93.213(-5.398)
Naive ViT-B/16 2.167(+0.364) 81.489(-12.983) 0.015(-1.788) 12.373(+6.342) 76.050(-22.640) 3.164(+1.361) 9.610(+3.579) 88.253(-10.437)
F3Net 5.743(+4.614) 87.867(-7.852) 3.521(+2.392) 6.445(+1.575) 85.112(-13.602) 1.188(+0.059) 16.306(+11.436) 91.603(-7.111)
SPSL 0.601(-0.436) 80.006(-16.340) 5.109(+4.072) 7.842(+3.308) 82.175(-16.572) 1.385(+0.348) 9.261(+4.272) 98.838(+0.091)
Frequency SRM 7.000(+5.228) 79.768(-15.324) 3.823(+2.051) 6.567(-0.213) 66.401(-31.535) 3.281(+1.509) 7.907(+1.127) 90.049(-7.887)
UCF 2.169(+0.718) 93.009(-2.142) 8.687(+7.236) 17.068(+11.558) 80.821(-17.261) 3.513(+2.062) 10.529(+5.019) 87.778(-10.304)
UnivFD 7.625(+6.003) 67.983(-25.668) 4.540(+2.918) 9.950(+4.542) 76.443(-21.749) 1.645(+0.023) 3.848(-1.560) 94.418(-3.774)
Spatial CORE 4.410(+3.021) 83.328(-12.896) 7.741(+6.352) 17.348(+11.417) 77.226(-21.353) 3.759(+2.370) 23.289(+17.358) 98.408(-0.171)
DAW-FDD 4.726(+3.401) 84.685(-10.741) 5.536(+4.211) 13.667(+7.791) 81.807(-16.004) 1.443(+0.118) 10.228(+4.532) 97.854(+0.043)
DAG-FDD 2.364(+0.944) 83.918(-11.804) 3.064(+1.644) 22.203(+16.137) 75.206(-23.565) 0.714(-0.706) 10.332(+4.266) 92.108(-6.663)
Fairness- enhanced PG-FDD 1.513(+0.442) 92.852(-3.322) 4.565(+3.494) 9.717(+6.015) 85.271(-13.901) 3.063(+1.992) 9.479(+5.777) 93.329(-5.843)
@@
Table 5: Fairness generalization results based on the gender attribute. The smallest performance changes (in parentheses) and the best performance are in red and in bold, respectively. Only $F_{OAE}$ fairness metric and ACC metric are used in CCv2 due to all samples are real.
@@
Figure 7: Impact of the training set size (Left) and the ratio of real and fake (Right) on detectors’ intersectional $F_{EO}$(%) and AUC (%).
Fairness Generalization Evaluation. To evaluate detectors’ fairness generalization capability, we test them on Casual Conversations v2 (CCv2) [99], DF-Platter [16], and GenData [17], none of which are part of AI-Face. Notably, CCv2 is a dataset that contains only real face images with demographic annotations (e.g., gender) self-reported by the participants. Results on gender attribute in Table 5 show that: 1) Even well-designed detectors that focus on improving utility or fairness generalization (e.g., UCF, PG-FDD) struggle to achieve consistently superior performance across different dataset domains. This highlights the remaining fairness generalization issue. 2) DAW-FDD and PG-PDD are two fairness-enhanced detectors that require accessing demographic information during training, but their fairness does not encounter a drastic drop when evaluating on CCv2. This reflects the high accuracy of the annotations in our AI-face.
Effect of Training Set Size. We randomly sample 20%, 40%, 60%, and 80% of each training subset from AI-Face to assess the impact of training size on performance. Key observations from Fig. 7 (Left): 1) Among all detectors, UnivFD demonstrates the most stable fairness and utility performance as the training dataset size changes, likely due to its fixed CLIP backbone. 2) Increasing the training dataset size generally improves model utility, but this pattern does not extend to fairness metrics. In fact, certain detectors such as F3Net and UCF exhibit worsening fairness as the training size reaches its maximum. This suggests that more training data does not necessarily lead to fairer detectors.
Effect of the Ratio of Real and Fake. To examine how training real-to-fake sample ratios affect detector performance, we set the ratios at 1:10, 1:1, and 10:1 while keeping the total sample count constant. Experimental results in Fig. 7 (Right) show: 1) Most detectors’ fairness improves as real sample representation increases. Probably because increasing real and reducing fake samples helps detectors reduce overfitting to artifacts specific to fake samples. This makes it easier for detectors to distinguish real from fake, even for underrepresented groups, thereby enhancing fairness. 2) Most detectors achieve the highest AUC with balanced data.
5.3 Discussion
@@
According to the above experiments, we summarize the unsolved fairness problems in recent detectors: 1) Detectors’ fairness is unstable when detecting face images generated by different generative methods, indicating a future direction for enhancing fairness stability since new generative models continue to emerge. 2) Even though fairness-enhanced detectors exhibit small overall fairness metrics, they still show biased detection towards minority groups. Future studies should be more cautious when designing fair detectors to ensure balanced performance across all demographic groups. 3) There is currently no reliable detector, as all detectors experience severe large performance degradation under image post-processing and cross-domain evaluation. Future studies should aim to develop a unified framework that ensures fairness, robustness, and generalization, as these three characteristics are essential for creating a reliable detector. Moreover, integrating foundation models (e.g., CLIP) into detector design may help mitigate bias.
@@ 6 Conclusion
This work presents the first demographically annotated million-scale AI-Face dataset, serving as a pivotal foundation for addressing the urgent need for developing fair AI face detectors. Based on this dataset, we conduct the first comprehensive fairness benchmark, shedding light on the fairness performance and challenges of current representative AI face detectors. Our findings can inspire and guide researchers in refining current models and exploring new methods to mitigate bias. Limitation and Future Work: One limitation is that our dataset’s annotations are algorithmically generated, so they may lack 100% accuracy. This challenge is difficult to resolve, as demographic attributes for most AI-generated faces are often too ambiguous to predict and do not map to real-world individuals. We plan to enhance annotation quality through human labeling in the future. We also plan to extend our fairness benchmark to evaluate large language models like LLaMA2 [100] and GPT4 [101] for detecting AI faces. Social Impact: Malicious users could misuse AI-generated face images from our dataset to create fake social media profiles and spread misinformation. To mitigate this risk, only users who submit a signed end-user license agreement will be granted access to our dataset.
Ethics Statement
Our dataset collection and annotation generation are approved by Purdue’s Institutional Review Board. The dataset is only for research purposes. All data included in this work are sourced from publicly available datasets, and we strictly comply with each dataset’s license agreement to ensure lawful inclusion and permissible secondary use for training and testing. All collected data and their associated licenses are mentioned in the Datasheet of AI-Face in Appendix E. Our annotation processes prioritize ethical considerations: 1) 76% images we annotated are generated facial images, ensuring no potential for harm to any individual. 2) For real images, we only provide annotations for content either licensed by the original copyright holders or explicitly stated as freely shareable for research purposes.
Acknowledgments
This work is supported by the U.S. National Science Foundation (NSF) under grant IIS-2434967 and the National Artificial Intelligence Research Resource (NAIRR) Pilot and TACC Lonestar6. The views, opinions and/or findings expressed are those of the author and should not be interpreted as representing the official views or policies of NSF and NAIRR Pilot.
@@ Acknowledgments
Appendix A The Definition of Demographic Categories
Skin Tone: Skin tone is an important attribute of human appearance, with significant variation from pale to dark. Recently, AI systems, especially computer vision models, have become controversial over concerns about the potential bias of performance varying based on skin tone [102, 46, 103]. Additionally, existing research has pointed out that skin tone annotations can be potentially less biased than building a racial classifier [104]. And the ethnicity attribute is subjective and can conceptually cause confusion in many aspects; for example, there may be no difference in facial appearance of African-American and African people, although, they may be referred to with two distinct racial categories. We, therefore, have opted to annotate the apparent skin tone of each face image. The Monk Skin Tone Scale [48] is developed specifically for the computer vision use case. We intentionally use the Monk Skin Tone scale over the Fitzpatrick skin type [105], which is developed as means for determining one’s likelihood of getting sunburn and lacks variance in darker skin tones [106, 107]. Additionally, Fitzpatrick skin type has been shown to be unreliable for image annotation [108].
Gender: Many governments [49, 50] have adopted binary gender (i.e., Man/Male (M) and Woman/Female(F), defined as sex at birth, as a common choice for legal and institutional systems and official documents. Most facial recognition research [45, 51, 52] also considers binary genders in their analyses. Our AI-Face dataset adopts binary gender as gender attributes.
Age: Follow United Nations [53] and Statistics Canada [54], we have five distinct perceived age groups- Child (0-14), Youth (15-24), Adults (25-44), Middle-age Adults (45-64), and Seniors (65+).
The demographic attribute and its corresponding example images are shown from Fig. A.1 to Fig. A.3.
Figure A.1: Demographic annotation definition and examples of skin tone attribute.
Figure A.2: Demographic annotation definition and examples of gender attribute.
Figure A.3: Demographic annotation definition and examples of age attribute.
Appendix B The Details of Demographically Annotated AI-Face Dataset
Figure B.1: Overview of AI-Face dataset. Each face has three demographic annotations.
B.1 Detailed Information of Datasets
We build our AI-Face dataset by collecting and integrating public real and AI-generated face images sourced from academic publications, GitHub repositories, and commercial tools. We strictly adhere to the license agreements of all datasets to ensure that they allow inclusion in our datasets and secondary use for training and testing. Table B.1 shows the detailed information of each dataset we used in our AI-Face, including the number of samples, the link for downloading the dataset, the accessibility, and their licenses.
Dataset #Samples Link Access License
FF++
 [2]
127K https://github.com/ondyari/FaceForensics/tree/master/dataset
freely shared for a research purpose,
submit aggreement
Non Commercial
DFDC
 [21]
75K https://www.kaggle.com/c/deepfake-detection-challenge/data Unknown
DFD
 [22]
40K https://research.google/blog/contributing-data-to-deepfake-detection-research/ the rights have been cleared for real videos, submit aggreement Non Commercial
Celeb-DF-v2
 [23]
179K https://cse.buffalo.edu/\$\$siweilyu/celeb-deepfakeforensics.html
freely shared for a research purpose,
submit aggreement
Non Commercial
AttGAN
 [55]
6K
StarGAN
 [55]
5.6K
StyleGAN
 [55]
10K https://iplab.dmi.unict.it/mfs/Deepfakes/PaperGANDCT-2021/ Online dataset, download directly, no license or agreement to sign Unknown
StyleGAN2
 [57]
118K https://github.com/SelfishGene/SFHQ-dataset
Since all images in this dataset are synthetically
generated there are no privacy issues or
license issues surrounding these images.
MIT License
StyleGAN3
 [58]
26.7K https://huggingface.co/datasets/InfImagine/FakeImageDataset
This dataset are fully open for academic research
and can be used for commercial purposes
with official written permission.
Apache-2.0
MMDGAN
 [56]
1K
MSGGAN
 [56]
1K
STGAN
 [56]
1K https://github.com/vishal3477/Reverse_Engineering_GMs/blob/main/dataset/ The dataset can be used for research purposes only and can be used for commercial purposes with official written permission. Non Commercial
ProGAN
 [59]
100K https://drive.google.com/drive/folders/1jU-hzyvDZNn_M3ucuvs9xxtJNc9bPLGJ
Online dataset, download directly,
no license or agreement to sign
Unknown
VQGAN
 [60]
50K https://github.com/awsaf49/artifact
This dataset comes from ArtiFact dataset,
which dataset takes leverage of data
from multiple methods thus different parts
of the dataset come with different licenses.
MIT License
DALLE2
 [61]
204
IF
 [61]
505
Midjourney
 [61]
100 https://github.com/ZhendongWang6/DIRE freely shared for a research purpose
DCFACe
 [62]
529K https://github.com/mk-minchul/dcface freely shared for a research purpose Unknown
Latent Diffusion
 [63]
20K https://github.com/grip-unina/DMimageDetection
Copyright 2024 Image Processing Research Group of
University Federico II of Naples (’GRIP-UNINA’).
All rights reserved.Licensed under the
Apache License, Version 2.0 (the ”License”)
Apache-2.0
Palette
 [64]
6K https://github.com/awsaf49/artifact/?tab=readme-ov-file#data-generation
This dataset comes from ArtiFact dataset,
which dataset takes leverage of data from
multiple methods thus different parts
of the dataset come with different licenses.
MIT License
SD v1.5
 [65]
18K
SD Inpainting
 [65]
20.9K https://huggingface.co/datasets/OpenRL/DeepFakeFace freely shared for a research purpose Apache-2.0
FFHQ
 [6]
70K https://github.com/NVlabs/ffhq-dataset
You can use, redistribute, and adapt it for
non-commercial purposes, as long as you
(a) give appropriate credit by citing our paper,
(b) indicate any changes that you’ve made,and
(c) distribute any derivative works under the same license.
Creative Commons
BY-NC-SA 4.0 license
IMDB-WIKI
 [20]
239K https://data.vision.ee.ethz.ch/cvl/rrothe/imdb-wiki/
This dataset is made available for academic
research purpose only. All the images are
collected from the Internet, and the
copyright belongs to the original owners.
Non Commericial
Table B.1: A list of datasets used in AI-Face, including the number of samples, links, access details, and licenses.
B.2 Artifacts of Deepfake Forgeries in Frequency
Leveraging frequency domain information plays a pivotal role in detecting AI-generated images. Frequency-based methods analyze the frequency components of an image, capturing information that may not be readily apparent in the spatial domain. In Fig. B.2, we present the mean Fast Fourier Transform (FFT) spectrum of images sampled from various sources in our AI-Face dataset. The results indicate that generative models often concentrate their output energy in the low-frequency range, represented by the central area of the FFT spectrum, resulting in overly smooth images. Notably, some models, such as StarGAN and Midjourney, exhibit distinct frequency artifacts, suggesting that they continue to struggle with eliminating generative patterns in the frequency domain. These artifacts serve as critical cues for distinguishing synthetic images from real ones. While most prior work has focused on applying frequency information to enhance the utility performance of detectors, exploring how frequency features can be leveraged to improve the fairness of detectors presents a promising direction for future research.
Figure B.2: Frequency analysis on various sources. The mean FFT spectrum computation involves averaging over 2,000 images. DALLE2, IF, and Midjourney take average over 200, 500, and 100, respectively, due to their small number of samples.
B.3 Experimental Study of Existing Face Attribute Prediction Tools
We compare current state-of-the-art face attribute prediction tools Face++ [71] and InsightFace [72] with our annotator. We perform intra-domain (train and test on IMDB-WIKI) and cross-domain (train on IMDB-WIKI, test on four AI-generated face datasets) evaluations. FF++, DFDC, DFD, and Celeb-DF-v2 are selected for cross-domain evaluation because they contain AI-generated faces, which match our objective and are not used to train Face++ and InsightFace. Additionally, they have demographic attribute annotations from [19], which can be used as ground truth for annotator evaluation. Since those annotations provided by  [19] have limited age annotations, our evaluation of these four datasets is confined to gender. The intra-domain results are shown in Table B.2 and the results of cross-domain are in Table B.3. Those results demonstrate our annotator’s superiority in demographic attribute prediction and generalization capability against Face++ and InsightFace. For example, under intra-domain evaluation (Table B.2), its precision surpasses the second-best method, InsightFace, by 3.47% on Female and 24.81% on Senior. In cross-domain evaluation (Table B.2), our annotator maintains high accuracy on all datasets, reflecting good generalization. For instance, on the DFDC dataset, the precision our annotator outperforms Face++ by a margin of up to 1.07% and InsightFace by 3.32% on Female.
Female Male Child Young Adult Mid Senior
Method Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
0.9161 0.9142 0.9151 0.9143 0.9163 0.9153 0.8579 0.0220 0.0429 0.3942 0.2940 0.3368 0.3215 0.6700 0.4345 0.5423 0.6460 0.5894 0.8078 0.7700 0.7884
Face ++ (0.0064) (0.0031) 0.0045 (0.0033) (0.0067) 0.0048 (0.0962) (0.0062) (0.0119) (0.0222) (0.0161) (0.0186) (0.0054) (0.0192) (0.0087) (0.0160) (0.0210) (0.0147) (0.0143) (0.0165) (0.0127)
0.9648 0.9405 0.9525 0.9420 0.9656 0.9537 1.0000 0.0020 0.0040 0.3970 0.0693 0.1180 0.2599 0.6180 0.3659 0.4105 0.5733 0.4783 0.7224 0.7560 0.7386
Insightface (0.0055) (0.0058) 0.0017 (0.0050) (0.0057) 0.0017 (0.0000) (0.0016) (0.0032) (0.0337) (0.0098) (0.0155) (0.0027) (0.0124) (0.0048) (0.0067) (0.0196) (0.0094) (0.0124) (0.0238) (0.0143)
0.9995 0.9992 0.9993 0.9992 0.9995 0.9993 0.9787 0.9780 0.9783 0.9560 0.9393 0.9476 0.9265 0.9547 0.9402 0.9498 0.9320 0.9408 0.9642 0.9700 0.9671
Ours (0.0006) (0.0006) 0.0006 (0.0006) (0.0006) 0.0006 (0.0027) (0.0062) (0.0039) (0.0078) (0.0080) (0.0042) (0.0122) (0.0107) (0.0058) (0.0123) (0.0123) (0.0124) (0.0085) (0.0087) (0.0082)